# Qwen3-ASR Technical Report

### v1

@@ 1 Introduction
In this report, we present the Qwen3-ASR family, including Qwen3-ASR-1.7B and Qwen3-ASR-0.6B - two all-in-one ASR models with language identification (LID) ability for 52 languages and dialects, and Qwen3-ForcedAligner-0.6B - the first lightweight LALM-based multilingual forced aligner supporting 11 languages and flexible timestamp prediction granularities. These model are posttrained from the strong foundation model of Qwen3-Omni (Xu et al., 2025b). For evaluating the performance of ASR models on benchmarks out of the open-sourced ones (ASR models at present have reached the limit of annotation errors on several test sets), we build a series of internal benchmarks covering more than complex acoustic environment, dialects, elders and kids speech and multilingual. Qwen3-ASR-1.7B achieves state-of-the-art (SOTA) performance among open-sourced ASR models and is competitive with the strongest proprietary commercial APIs. Qwen3-ASR-0.6B offers the best accuracy-model-size trade-off, making it a strong choice for on-device deployment. Qwen3-ForcedAligner-0.6B delivers highly accurate forced-alignment timestamps and inherits the key capabilities of Qwen3-ASR, including multilingual and long-form speech support, enabling scalable labeling of speech-transcript pairs.
@@
  • Presents novel speech force alignment architecture. To the best of our knowledge, we introduce the first Large Language Model based speech forced aligner that produces accurate timestamps at flexible granularities, including word, sentence, and paragraph levels. Compared with forced-alignment models at present including Montreal Forced Aligner (MFA) and NeMo Forced Aligner (NFA), Qwen3-ForcedAligner-0.6B fills the gap of an all-in-one, multilingual speech forced alignment solution, and completes an essential function of the Qwen3-ASR family.
  • Open-sourced weights, powerful inference and finetune framework. Except weights of the 3 released models, we also open-source complete and easy-using code base for inference with multiple features as well as a finetuning recipe. We hope this unified toolkit facilitates the research and development of the ASR community.
@@ 2.1 Architecture
Qwen3-ASR famliy models leverage Qwen3-Omni as foundation model, which proved to obtain strong audio understanding ability Xu et al. (2025a). Speech to recognize is first fed to AuT encoder, which is pretrained separately from Qwen3-Omni and Qwen3-ASR. As illustrated in Figure 2(left), AuT is an attention-encoder-decoder (AED) based ASR model which conducts 8 times downsampling to Fbank feature with 128 dimensions, yielding a 12.5Hz token rate audio encoder. We use dynamic flash attention window size ranging from 1s to 8s, which allows Qwen3-ASR to perform both streaming inference with short chunks and offline inference with long queries. The architecture of models we are releasing is illustrated as Figure 2(right) and detailed as below: Qwen3-ASR-1.7B is built with Qwen3-1.7B, a projector and the an AuT encoder with 300M parameters and 1024 hidden size. This model demonstrates strong performance in multilingual and dialect speech recognition, so as to robustness under complex acoustic environment and text patterns. Qwen3-ASR-0.6B is built with Qwen3-0.6B, a projector and an AuT encoder with 180M parameters and a hidden size of 896. We design this compact model to balance recognition accuracy and inference efficiency, while remaining highly competitive among sub-1B-parameter ASR models.
@@ 2.2 Training Strategies
    ASR supervised finetuning (SFT). In the SFT stage, we perform style transfer on the ASR input/output format with a substantially smaller set of multilingual data that is disjoint from the pretraining corpus. Besides standard Chinese, English and multilingual ASR data, SFT stage also utilize non-speech data, streaming-enhancement data and context biasing data. Specifically, we train the model to be an ASR-only model that does not follow natural-language instructions in the prompt, in order to mitigate instruction injection and instruction-following failures. The output of Qwen3-ASR has outputs in two types for the given audio with and without recognizable human speech:
@@
    Meanwhile, the model learns to utilize the context tokens inside the system prompt as background knowledge, allowing user to obtain customized ASR results.
@@
    ASR reinforcement learning (RL). At the last stage, we use Group Sequence Policy Optimization (GSPO, Zheng et al. (2025)) for further improving the quality of recognition. It turns out that RL plays an essential role for models’ noise robustness, transcribing stability and ability to analyze difficult cases. The total data leveraged by RL stage is about 50k utterances including 35% Chinese and English data, 35% multilingual data and 30% functional data which aims at improving transcribing stability in complex environments.
@@ 2.3 Features
With architecture and training strategies introduced above, Qwen3-ASR family models are notable in aspects below:
@@
    Long-form and streaming inference. Qwen3-ASR-1.7B and Qwen3-ASR-0.6B naturally supports single speech no longer than 20 minutes and streaming/offline unified inference.
@@ 2.4 Inference Efficiency
The speed benchmarks of Qwen3-ASR are conducted in two settings: offline batch inference and online asynchronous inference. The former is evaluated using vLLM’s offline batch generation, while the latter is evaluated with a multi-concurrency request setup based on vLLM Serve, which better reflects inference efficiency in industrial environments. All experiments are run with vLLM v0.14.0, with CUDA Graph enabled and bfloat16 precision for inference. The results in Table 2 show that, under different concurrency levels, Qwen3-ASR-0.6B can achieve an average Time-to-First-Token (TTFT) as low as 92ms. It reaches real-time factor (RTF) as low as 0.064 and throughput as high as 2000 at a concurrency of 128, which means it can process 2,000 seconds of audio per second.
Table 2: Efficiency of Qwen3-ASR family models. Qwen3-ASR-0.6B and Qwen3-ASR-1.7B support vLLM based inference in both offline batch and online asynchronous mode, while Qwen3-ForcedAligner-0.6B supports offline batch inference in PyTorch only. All measurements in the table are based on input audio of approximately 2 minutes for ASR and 1 minutes for FA in length, and all inference is performed on a single typical compute resource. Conc. denotes the concurrency level. TTFT p95 denotes the 95th percentile TTFT latency.
@@ 3.1 Overview
  • Broad Application Scenarios. Qwen3-ForcedAligner-0.6B supports speech in 11 languages with durations of up to 300 seconds, including the cross-lingual scenarios, and allows users to flexibly customize timestamp prediction for any word or character.
@@ 3.2 Model Design
Figure 3: Illustration of Qwen3-ForcedAligner-0.6B. During training, randomly masked timestamp slots with are dynamically inserted into the token sequence to represent word or character boundaries. The combined sequence is fed into Qwen3-0.6B LLM, and a timestamp prediction layer predicts the corresponding timestamp indices for each slot. Supervision is applied with cross‑entropy loss on synchronously aligned label and output sequences.
@@ 4.2.1 Opensource ASR Benchmarks
As shown in Table A.1, Qwen3-ASR delivers consistently strong performance across English, Mandarin Chinese, and multiple Chinese dialect benchmarks. It is competitive with leading commercial APIs while substantially outperforming widely used open-source baselines. Scaling from Qwen3-ASR-0.6B to Qwen3-ASR-1.7B yields clear and stable gains, indicating that the model benefits effectively from increased capacity.
@@
Overall, Table A.1 highlights three key advantages of Qwen3-ASR: (i) strong cross-domain generalization on English benchmarks beyond curated read speech, (ii) state-of-the-art accuracy on Mandarin Chinese across multiple public datasets including large-scale, noisy meeting-style speech, and (iii) robust handling of Chinese dialects, with especially strong performance on Cantonese and long/short dialectal speech. These findings demonstrate that Qwen3-ASR delivers strong, reproducible performance across diverse public benchmarks, while also remaining competitive with top-tier closed-source APIs.
Table 3: Evaluation on English, Mandarin Chinese, and a range of Chinese dialect benchmarks. For the commercial APIs and the open-source Whisper-large-v3 model, we obtained results by running inference on the test sets ourselves due to the absence of published numbers; for FunASR-MLT-Nano, we report the results from its official technical report. ”N/A” denotes we cannot get a reasonable result by the official API. ”–” indicates that the corresponding benchmark result is not reported.
@@ 4.3.1 Multilingual ASR Performance
On MLS, Common Voice and MLC-SLM, Qwen3-ASR-1.7B consistently outperforms the evaluated open-source baselines, including the widely used Whisper-large-v3, and substantially surpasses smaller multilingual models. On Fleurs, which spans more languages and diverse recording conditions, Qwen3-ASR-1.7B achieves the best performance on the 12- and 20-language subsets. However, relative to Whisper-large-v3, its performance degrades on the full 30-language setting, indicating room for improvement in handling increased linguistic diversity and long-tail languages. Nevertheless, Qwen3-ASR-1.7B remains markedly better than the 0.6B variant, suggesting that model scaling improves robustness in more challenging multilingual regimes.
@@ 4.3.2 Language Identification Performance
Following the output template in Section 2.2, Qwen3-ASR not only decodes speech into text, but also performs language identification (LID) via natural-language prompting before ASR decoding. In this section, we evaluate LID accuracy on 4 multilingual benchmarks: Fleurs (30 languages), MLS (9 languages), CommonVoice (13 languages), MLC-SLM (11 languages); the covered languages are detailed in Section 2.3. As shown in Table 6, we compare Qwen3-ASR-0.6B and Qwen3-ASR-1.7B with Whisper-large-v3, a strong multilingual ASR model with built-in LID capability. Both Qwen3-ASR models outperform Whisper-large-v3, demonstrating stable and effective language identification across these mainstream languages. Most remaining errors on Fleurs stem from confusions between Malay (ms) and Indonesian (id), two closely related languages with high acoustic similarity.
@@ 5 Conclusion
In this paper we introduce Qwen3-ASR model family, including two ASR models and a forced-alignment model trained with large-scale speech data. By leveraging the strong audio understanding ability of foundation model Qwen3-Omni and our training process with 4 stages, Qwen3-ASR-1.7B and Qwen3-ASR-0.6B outperform competing models of similar or larger size and commercial APIs in both speech coverage and recognition accuracy. They support language identification and speech recognition for 30 languages, conduct accurate ASR under complex acoustic environment, maintain robustness for accents and dialects, and keep performant on singing voice and other speech from real-world scenarios. Besides ASR models, we also propose a novel solution for timestamp prediction and forced alignment: Qwen3-ForcedAligner-0.6B, which is an LLM based NAR timestamp predictor that supports FA for 11 languages and within 5 minutes. It outperforms three mainstream end-to-end ASR model based FA solutions in timestamp accuracy, inference speed and language coverage. Together with the weights of three models, we also open-source a powerful and easy-to-use inference framework. Overall, the Qwen3-ASR family achieves state-of-the-art performance on both real-world evaluations and public benchmarks, and our open-sourced forced-alignment model completes a critical missing piece in the speech technology stack. We hope Qwen3-ASR can contribute to the advancement of speech recognition research and applications, and we will continue to improve our open model family in both accuracy and functional innovations.

### v2

@@ 1 Introduction
In this report, we present the Qwen3-ASR family, including Qwen3-ASR-1.7B and Qwen3-ASR-0.6B - two all-in-one ASR models with language identification (LID) ability for 52 languages and dialects, and Qwen3-ForcedAligner-0.6B - the first lightweight LALM-based multilingual forced aligner supporting 11 languages and flexible timestamp prediction granularities. These models are posttrained from the strong foundation model of Qwen3-Omni (Xu et al., 2025b). For evaluating the performance of ASR models on benchmarks out of the open-sourced ones (ASR models at present have reached the limit of annotation errors on several test sets), we build a series of internal benchmarks covering more than complex acoustic environment, dialects, elders and kids speech and multilingual. Qwen3-ASR-1.7B achieves state-of-the-art (SOTA) performance among open-sourced ASR models and is competitive with the strongest proprietary commercial APIs. Qwen3-ASR-0.6B offers the best accuracy-model-size trade-off, making it a strong choice for on-device deployment. Qwen3-ForcedAligner-0.6B delivers highly accurate forced-alignment timestamps and inherits the key capabilities of Qwen3-ASR, including multilingual and long-form speech support, enabling scalable labeling of speech-transcript pairs.
@@
  • Presents novel speech force alignment architecture. To the best of our knowledge, we introduce the first Large Language Model based speech forced aligner that produces accurate timestamps at flexible granularities, including word, sentence, and paragraph levels. In contrast to existing tools such as the Montreal Forced Aligner (MFA) and NeMo Forced Aligner (NFA), our model, Qwen3-ForcedAligner-0.6B, offers a unified, multilingual solution that addresses the lack of an all-in-one forced alignment system within the Qwen3-ASR family and fulfills a critical functional component for comprehensive spoken language processing.
  • Open-source models and a comprehensive inference and fine-tuning framework. In addition to releasing the weights of three models, we provide a fully open-source, user-friendly codebase that supports inference with multiple features (e.g., multi-granularity alignment, streaming transcription, and multilingual processing) as well as a reproducible fine-tuning recipe. We hope this unified toolkit will accelerate research and development efforts in the automatic speech recognition community.
@@ 2.1 Architecture
Qwen3-ASR family models leverage Qwen3-Omni as a foundation model, which proved to obtain strong audio understanding ability (Xu et al., 2025a). Speech to recognize is first fed to AuT encoder, which is pretrained separately from Qwen3-Omni and Qwen3-ASR. As illustrated in Figure 2(left), AuT is an attention-encoder-decoder (AED) based ASR model which conducts 8 times downsampling to Fbank feature with 128 dimensions, yielding a 12.5Hz token rate audio encoder. We use a dynamic flash attention window size ranging from 1s to 8s, which allows Qwen3-ASR to perform both streaming inference with short chunks and offline inference with long queries. The architecture of models we are releasing is illustrated as Figure 2(right) and detailed as below: Qwen3-ASR-1.7B is built with Qwen3-1.7B, a projector, and an AuT encoder with 300M parameters and 1024 hidden size. This model demonstrates strong performance in multilingual and dialect speech recognition, so as well as robustness under complex acoustic environment and text patterns. Qwen3-ASR-0.6B is built with Qwen3-0.6B, a projector and an AuT encoder with 180M parameters and a hidden size of 896. We design this compact model to balance recognition accuracy and inference efficiency, while remaining highly competitive among sub-1B-parameter ASR models.
@@ 2.2 Training Strategies
    ASR supervised finetuning (SFT). In the SFT stage, we perform style transfer on the ASR input/output format with a substantially smaller set of multilingual data that is disjoint from the pretraining corpus. Besides standard Chinese, English and multilingual ASR data, SFT stage also utilizes non-speech data, streaming-enhancement data and context biasing data. Specifically, we train the model to be an ASR-only model that does not follow natural-language instructions in the prompt, in order to mitigate instruction injection and instruction-following failures. The output of Qwen3-ASR has outputs in two types for the given audio, with and without recognizable human speech:
@@
    Meanwhile, the model learns to utilize the context tokens inside the system prompt as background knowledge, allowing users to obtain customized ASR results.
@@
    ASR reinforcement learning (RL). At the last stage, we use Group Sequence Policy Optimization (GSPO, Zheng et al. (2025)) for further improving the quality of recognition. It turns out that RL plays an essential role in models’ noise robustness, transcription stability and ability to analyze difficult cases. The total data leveraged by RL stage is about 50k utterances including 35% Chinese and English data, 35% multilingual data and 30% functional data, which aims at improving transcribing stability in complex environments.
@@ 2.3 Features
With the architecture and training strategies introduced above, Qwen3-ASR family models are notable in the aspects below:
@@
    Long-form and streaming inference. Qwen3-ASR-1.7B and Qwen3-ASR-0.6B naturally support single speech no longer than 20 minutes and streaming/offline unified inference.
@@ 2.4 Inference Efficiency
The speed benchmarks of Qwen3-ASR are conducted in two settings: offline batch inference and online asynchronous inference. The former is evaluated using vLLM’s offline batch generation, while the latter is evaluated with a multi-concurrency request setup based on vLLM Serve, which better reflects inference efficiency in industrial environments. All experiments are run with vLLM v0.14.0, with CUDA Graph enabled and bfloat16 precision for inference. The results in Table 2 show that, under different concurrency levels, Qwen3-ASR-0.6B can achieve an average Time-to-First-Token (TTFT) as low as 92ms. It reaches a real-time factor (RTF) as low as 0.064 and throughput as high as 2000 at a concurrency of 128, which means it can process 2,000 seconds of audio per second.
Table 2: Efficiency of Qwen3-ASR family models. Qwen3-ASR-0.6B and Qwen3-ASR-1.7B support vLLM-based inference in both offline batch and online asynchronous mode, while Qwen3-ForcedAligner-0.6B supports offline batch inference in PyTorch only. All measurements in the table are based on input audio of approximately 2 minutes for ASR and 1 minute for FA in length, and all inference is performed on a single typical computing resource. Conc. denotes the concurrency level. TTFT p95 denotes the 95th percentile TTFT latency.
@@ 3.1 Overview
  • Broad Application Scenarios. Qwen3-ForcedAligner-0.6B supports speech in 11 languages with durations of up to 300 seconds, including cross-lingual scenarios, and allows users to flexibly customize timestamp prediction for any word or character.
@@ 3.2 Model Design
Figure 3: Illustration of Qwen3-ForcedAligner-0.6B. During training, randomly masked timestamp slots are dynamically inserted into the token sequence to represent word or character boundaries. The combined sequence is fed into Qwen3-0.6B LLM, and a timestamp prediction layer predicts the corresponding timestamp indices for each slot. Supervision is applied with cross‑entropy loss on synchronously aligned label and output sequences.
@@ 4.2.1 Opensource ASR Benchmarks
As shown in Table 3, Qwen3-ASR delivers consistently strong performance across English, Mandarin Chinese, and multiple Chinese dialect benchmarks. It is competitive with leading commercial APIs while substantially outperforming widely used open-source baselines. Scaling from Qwen3-ASR-0.6B to Qwen3-ASR-1.7B yields clear and stable gains, indicating that the model benefits effectively from increased capacity.
@@
Overall, Table 3 highlights three key advantages of Qwen3-ASR: (i) strong cross-domain generalization on English benchmarks beyond curated read speech, (ii) state-of-the-art accuracy on Mandarin Chinese across multiple public datasets including large-scale, noisy meeting-style speech, and (iii) robust handling of Chinese dialects, with especially strong performance on Cantonese and long/short dialectal speech. These findings demonstrate that Qwen3-ASR delivers strong, reproducible performance across diverse public benchmarks, while also remaining competitive with top-tier closed-source APIs.
Table 3: Evaluation on English, Mandarin Chinese, and a range of Chinese dialect benchmarks. For the commercial APIs and the open-source Whisper-large-v3 model, we obtained results by running inference on the test sets ourselves due to the absence of published numbers; for FunASR-MLT-Nano, we report the results from its official technical report. ”N/A” denotes that we cannot get a reasonable result by the official API. ”–” indicates that the corresponding benchmark result is not reported.
@@ 4.3.1 Multilingual ASR Performance
On MLS, Common Voice and MLC-SLM benchmarks, Qwen3-ASR-1.7B consistently outperforms the evaluated open-source baselines, including the widely used Whisper-large-v3, and substantially surpasses smaller multilingual models. For Fleurs, which spans more languages and diverse recording conditions, Qwen3-ASR-1.7B achieves the best performance on the 12- and 20-language subsets. However, relative to Whisper-large-v3, its performance degrades on the full 30-language setting, indicating room for improvement in handling increased linguistic diversity and long-tail languages. Nevertheless, Qwen3-ASR-1.7B remains markedly better than the 0.6B variant, suggesting that model scaling improves robustness in more challenging multilingual regimes.
@@ 4.3.2 Language Identification Performance
Following the output template in Section 2.2, Qwen3-ASR not only decodes speech into text, but also performs language identification (LID) via natural-language prompting before ASR decoding. In this section, we evaluate LID accuracy on 4 multilingual benchmarks: Fleurs (30 languages), MLS (9 languages), CommonVoice (13 languages), MLC-SLM (11 languages); the covered languages are detailed in Section 2.3. As shown in Table 6, we compare Qwen3-ASR-0.6B and Qwen3-ASR-1.7B with Whisper-large-v3, a strong multilingual ASR model with built-in LID capability. Both Qwen3-ASR models outperform Whisper-large-v3, demonstrating stable and effective language identification across these mainstream languages. Most remaining errors on Fleurs stem from confusion between Malay (ms) and Indonesian (id), two closely related languages with high acoustic similarity.
@@ 5 Conclusion
We present Qwen3-ASR, a model family comprising two automatic speech recognition (ASR) systems and a forced-alignment (FA) model trained on large-scale speech corpora. By leveraging the strong audio understanding capability of the foundation model Qwen3-Omni and a four-stage training pipeline, Qwen3-ASR-1.7B and Qwen3-ASR-0.6B outperform competing models of comparable or larger size, as well as commercial APIs, in both speech coverage and recognition accuracy. The models support language identification and ASR across 30 languages, deliver robust performance in complex acoustic conditions, exhibit resilience to accents and dialects, and maintain effectiveness on singing voice and other real-world speech scenarios. In addition, we introduce Qwen3-ForcedAligner-0.6B, an LLM-based non-autoregressive timestamp predictor that enables forced alignment for 11 languages with end-to-end processing times under five minutes. This approach surpasses three mainstream end-to-end ASR-based FA solutions in timestamp accuracy, inference speed, and language coverage. Alongside releasing the weights for all three models, we open-source a unified and user-friendly inference framework. Overall, the Qwen3-ASR family achieves state-of-the-art performance on real-world evaluations and public benchmarks, and the open-sourced forced-alignment model addresses a critical gap in the speech technology stack. We will continue to advance this open model family in accuracy and functionality.